OpenStudy (anonymous):

for every positive 2 digit number, x, with tens digit t and units digit u, let y be the 2 digit number formed by reversing the digits of x. which of the following expressions is equivalent to x-y? f. 9(t-u) g. 9(u-t) h. 9t-u j. 9u-t k. 0

OpenStudy (anonymous):

i looked this one up...I know the answer, but i honestly can't figure out how they got it...

OpenStudy (anonymous):

lol yeah i dont know how you would get it either....

OpenStudy (anonymous):

i could look in the back of the book for the answer but i try to figure them out first but i have no flippin idea even where to start with this one

OpenStudy (anonymous):

x=10t+U y=10u+t x-y=10t+u-(10u+t) x-y=9t-9u=9(t-u)

OpenStudy (anonymous):

i understand everything up until the last line...I don't know where they get the nines from

OpenStudy (anonymous):

They just collected the terms of t and u

OpenStudy (anonymous):

how do you mean

OpenStudy (anonymous):

i do its called the foil method... hang on i will draw it out

OpenStudy (anonymous):

Looking at t: 10t -(+t) = 9t etc etc

OpenStudy (anonymous):

got ya...wow...ok over thought that

OpenStudy (anonymous):

in which he just showed you its the foil method

OpenStudy (anonymous):

ok you want another one?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

lol ok i will open it in a new window hun.

OpenStudy (anonymous):

alright babe

OpenStudy (anonymous):

i have to draw it out so i will just randomly type something..

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

x = 10t + u y = 10u + t x - y =10t + u - (10u +t) = 9t - 9u = 9(t-u)

OpenStudy (anonymous):

f is the answer